INAF queue observing, 2014 Oct 31-Nov 1 Halloween

Observer: E. Sani, R. Carini
Telescope Operator: S. Allanson
Support Astronomer: M. Ewards on call


House keeping problems for LBCBlue. Fail saving mode that stops observations. Is anyhow offered for science
00:36 is not possible open the dome, the temperature is too near the dew point and the humidity is 99.9%, it's raining on the horizon.
04.35 dome closed, strong wind, the temperature reaches the dew point.
09:45 dwepoint drops down, but still wind gusts are above 20 m/s.


05:20 we run the mods calibrations fos Israel A1 project.
  • execMODS catsj021950_ls_flat.cal mods1r.20141101.0002-0006, mods1b.20141101.0002-0006 flat programma Israel
  • execMODS slitflats_5as.cal mods1r.0006-0009 flat RedGrating slit 5"
  • execMODS grlamps_5arc.cal mods1b.0007-0009, mods1r.0009-0012 wrong slit width: 0.6", low counts
  • execMODS slitflats_5arc.cal mods1r.0013-0021, mods1b.0010-0018 flat program Israel
  • execMODS dual_grating_0.8_flat.cal mods1r.0022-0027, mods1b.0019-0024 flat for program Antoniucci
  • execMODS dual_grating_0.8_arc.cal mods1r.0027-0033, mods1b.0025-0035 arc for program Antoniucci

-- %USERSIG{RobertaCarini - 2014-11-01}%


Topic revision: r7 - 02 Nov 2014, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback